OpenStudy (anonymous):

During 2009, a particular country's inflation rate averaged 1.3% a day. This means that on average, prices went up by about 1.3% from one day to the next. Round your answer to the nearest whole number. (a) By what percentage did this country's prices increase in April of 2009 (from April 1 through April 30)?

OpenStudy (anonymous):

29 days of 1.3% inflation. Convert to relative increase: (1+0.013) = 1.013. (1.013)^29 -1 = 45.43% (the effect of compounding) 45%.

OpenStudy (anonymous):

Do you agree?

OpenStudy (anonymous):

yes thanks its more like 47 percent

OpenStudy (anonymous):

theres 30 days not 29